OpenStudy (anonymous):

The Tools Plus Company is planning to produce drills. The company has rented space for its manufacturing operation at $4,000 per month. Each drill requires $44 worth of material and $36 worth of labor. Find the total cost if Tools Plus Company makes 330 drills.

OpenStudy (bahrom7893):

It depends on how long it will take for them to make one drill.. not enough info

OpenStudy (bahrom7893):

If we don't count the rent: 330*(44+36) = 330*80=2640

OpenStudy (anonymous):

wrong. :(

OpenStudy (anonymous):

4000 + 330(44+36) = 6640

OpenStudy (a_clan):

2640+4000= $6640 per month